OpenStudy (anonymous):

x^2-36 over x^2-12+36

OpenStudy (anonymous):

What are you looking for? the answer of the factors

OpenStudy (anonymous):

yes @SebHuebl

OpenStudy (anonymous):

\[x^2-36 \over x^2+24\]

OpenStudy (anonymous):

thats not an option

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

If you happen to come by these types of questions they're really easy to solve especially since theirs a website called "mathway.com" just plug in your equation and it'll solve it for you.

OpenStudy (nurali):

\[\frac{ x^2-36 }{ x^2-12x+36 }\] \[\frac{ (x-6)(x+6) }{ (x-6)(x-6) }\] \[\frac{ (x+6) }{ (x-6) }\]

OpenStudy (anonymous):

The roots are 6 and -6

OpenStudy (anonymous):

Nurali has the work

OpenStudy (anonymous):

thank you all ! :D

OpenStudy (nurali):

Anytime.